OpenStudy (jhannybean):

Calculus 3: Trying to understand how to change from the xy-pane to the uv-plane to find the equation for a transformation where sides of S are parallel to u- and v- axes: I'm given \(y=3x-2\) , \(y=3x+4\) \ , \(y=-\frac{1}{2}x+1\) \ , \(y= -\frac{1}{2}x+3\)

OpenStudy (jhannybean):

@dan815

OpenStudy (dan815):

sounds like change of basis problem

OpenStudy (dan815):

jacobian stuff comes just out of partial derivaives and the cross product of the PD vectors

OpenStudy (jhannybean):

yeah... those are what I'm given :(

OpenStudy (jhannybean):

I'm not quite understanding how to make a change.

OpenStudy (dan815):

okay so

OpenStudy (dan815):

u see how its linear changes wrt to the intercepts

OpenStudy (dan815):



OpenStudy (dan815):

y=3x-2 y-3x=b, b from -2 to 4 y=-x/2+1 y+x/2=c, c from 1 to 3

OpenStudy (dan815):